Using the ideal gas law , calculate the pressure of O2 (g) in atm and the cylinder at 296 k.
1
Expert's answer
2018-05-18T11:16:08-0400
The partial pressure of oxygen in the atmospheric air is 20.94% of 760 mm, i.e. 159 mm(p1)
T1 = 273 K
T2 = 296 K
pV = nRT
if V(O2) = const, n(O2) = const =>
=> p1/p2 = T1/T2
p2 = p1*T2 / T1 = 159 mm*296 K/ 273K = 172.39 mm = 0.226 atm.